OpenStudy (anonymous):

Find the equation of the quadratic that has x-intercepts f(2)=0 and f(-6)=0 in vertex form.

OpenStudy (amistre64):

vertex x is half way between intercepts

OpenStudy (amistre64):

so its at least got vertex on the x=-2 axis

OpenStudy (anonymous):

how would i set it up then

OpenStudy (amistre64):

2+-6 = -4/2 = -2

OpenStudy (anonymous):

then how would I find the y

OpenStudy (amistre64):

f(x) = a(x+2)^2 + h

OpenStudy (amistre64):

err, i think its +k for standard usage

OpenStudy (anonymous):

yeah but how would i figure out k and the slope?

OpenStudy (amistre64):

without any other information; your pretty much at a stand still; there is a whole family of equations to fit this

OpenStudy (amistre64):

if k is negative; make a + and f(x) = a(x+2)^2 - k or the other way around since the vertex needs to be on the oterside of the opening end f(x) -a(x+2)^2 + k

OpenStudy (amistre64):



OpenStudy (amistre64):

anything but zeros for k and a should work, and make sure they gots opposite signs

OpenStudy (anonymous):

one sec let me test it

OpenStudy (anonymous):

but whats the slope then?

OpenStudy (amistre64):

hmmm, i just tested a few on the wolf, my thoughts need some refinement :)

OpenStudy (amistre64):

ok, try this; make k equal to half the distance between intercepts; in this case -4 should work for i trial run

OpenStudy (amistre64):

.... recalculating lol we are sitting at 4; so if x=4 x^2 = 16 .... lets try -16 hope this aint to confusing yet

OpenStudy (amistre64):

that one hit, i knew one of my ideas had to stick ;)

OpenStudy (anonymous):

so the slope is 1?

OpenStudy (amistre64):



OpenStudy (amistre64):

yes